# Data Trade and Consumer Privacy

### v2

@@ 4.1 Intermediary Optimality for $\mu_{0}>v_{L}/v_{H}$
When the mismatch cost in service is low as a penalty for the type-H consumers’ misreport $\delta^{2}<\Delta v$, the mismatch itself cannot provide sufficient incentive for the type-H consumers to reveal their information. Instead, the intermediary would like to lower $\pi\left(l|L\right)$ to incentivize truth-telling to collect information. In this case, some type-L consumers are excluded from the product market. The intermediary’s optimal profit from the service market is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and profit from the data trade is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. The equilibrium total payoff to the intermediary is $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$ and social surplus is $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}+\mu_{0}v_{H}$.
Following Proposition 1, we have the following results.
Corollary 1.
Suppose $\mu_{0}>v_{L}/v_{H}$. No information data trade cannot be optimal.
@@
To see it by contradiction. Suppose in the optimal mechanism $\pi\left(l|H\right)=\pi\left(l|L\right)$. In either case of $\delta^{2}\geq\Delta v$ or $\delta^{2}<\Delta v$, the obedience constraint (9) requires $\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}<\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)$ which contradicts with $\mu_{0}>\frac{v_{L}}{v_{H}}$. Intuitively, when $\mu_{0}>\frac{v_{L}}{v_{H}}$ any extra information will make it easier for the producer to price discriminate the consumers. This is because in the case of $\mu_{0}>\frac{v_{L}}{v_{H}}$ the market inclusion effect is positive: without any data trade all the type-L consumers are excluded in the product market.
Corollary 1 works as a rationale for the so-called digital “privacy paradox”: the general discrepancies between consumers’ active data-sharing behaviors and their self-stated privacy concerns. When $\mu_{0}>\frac{v_{L}}{v_{H}}$, trading some (even full) data is optimal. In other words, the highest privacy protection is not preferred. The results rationalize the paradox as a market outcome. Consumers would prefer to reveal some (even full) information about their private preferences even though they are informed that their revealed information might be traded and used to price against them in other markets. This is possible when the benefits of better service outweigh the cost of losing surpluses in the product. So the intermediary does not need to reduce probability $\pi\left(l|L\right)$. The choice of $\pi\left(l|H\right)$ also makes the producer obedient.
4.2 Intermediary Optimality for $\mu_{0}<v_{L}/v_{H}$
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, $\max\{v_{H}-p_{0},0\}=\Delta v$ and the obedience constraint for the producer to charge a low price in (9) is redundant. The participation constraint of the type-H consumers (13) is [^21] In the special parameter case of $\mu_{0}=v_{L}/v_{H}$, $p_{0}$ equals to $v_{H}$ or $v_{L}$ and then (13) is either (17) or (18).
Without data trade the producer will charge $v_{L}$ and all consumers purchase in the product market. The producer’s revenue is $R_{0}=v_{L}$ without data trade. The net gain in the product revenue from data trade (11) becomes
Here, the market inclusion effect of data trade is negative: With data trade, the market serving the type-L consumers shrinks from $1$ to $\pi\left(l|L\right)$. The surplus extraction effect of data trade is positive: The intermediary extracts extra surplus from the type-H consumers by $\left[1-\pi\left(l|H\right)\right]\left(v_{H}-v_{L}\right)$. As we will see, the extra extracted surplus due to data trade will be shifted back to the consumers in the service market.
Proposition 2.
Suppose $\mu_{0}<v_{L}/v_{H}$. In the optimal mechanism, the intermediary sets service fees to extract all the consumers surplus in the service and product markets
$f_{L}=V,\text{ }f_{H}=V-\left(1-\pi\left(l|H\right)\right)\Delta v$
and the optimal data trade depends on the cross-market comparison between the mismatch cost $\delta^{2}$ and the vertical differentiation $\Delta v$
$\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\left[\max\left\{1-\frac{\delta^{2}}{\Delta v},0\right\},1\right].$
The data trade fee
$T=\mu_{0}\Delta v\left(1-\pi\left(l|H\right)\right).$
and social welfare is
$V+v_{L}+\mu_{0}\Delta v.$
Similarly, $\pi\left(l|H\right)$ is canceled out in the intermediary’s objective due to the offset of surplus-shifting in the product and surplus-extracting in the service. By plugging the binding $IR_{L}$ (15) and $IR_{H}$ (18): $f_{L}=V$, $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$ into the objective, the intermediary solves the reduced problem subject to $IC_{L}$ (14) and $OB_{H}$ (8)
$\max_{\pi}$ $V+\left(1-\mu_{0}\right)v_{L}\pi\left(l|L\right)-\left(1-\mu_{0}\right)v_{L}$
subject to $\left(1-\pi\left(l|H\right)\right)\Delta v-\delta^{2}\leq 0$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\geq\frac{v_{L}}{v_{H}}-\mu_{0}$
The intermediary would set $\pi\left(l|L\right)=1$ as there is no constraint on it. At $\pi\left(l|L\right)=1$, any $\pi\left(l|H\right)\in\left[0,1\right]$ satisfies $OB_{H}$ (8). But $IC_{L}$ (14) imposes the restrictions on the choice set of $\pi\left(l|H\right)$, which relies on the mismatch as a penalty for misreporting. When the mismatch is high $\delta^{2}>\Delta v$, the penalty of service mismatch itself can incentivize type-L consumers to report the truth. So the intermediary can choose any $1-\pi\left(l|H\right)\in\left[0,1\right]$. In this case, the service profit is $V-\mu_{0}\Delta v\left[1-\pi\left(l|H\right)\right]$ and the data trade profit is $\mu_{0}\Delta v\left[1-\pi\left(l|H\right)\right]$. So its optimal total payoff is $V$.
When the mismatch is low $\delta^{2}<\Delta v$, the penalty of service mismatch itself cannot provide strong incentives for type-L consumers to tell the truth. The intermediary chooses $1-\pi\left(l|H\right)\in\big{[}0,\frac{\delta^{2}}{\Delta v}\big{]}$ such that the truth is induced. In this case, the optimal service profit is $V-\mu_{0}\left(1-\pi\left(l|H\right)\right)\Delta v$ and the data trade profit is $\mu_{0}\left(1-\pi\left(l|H\right)\right)\Delta v$. The intermediary’s total payoff is $V$.
In both cases, social welfare is the sum of intermediary surplus $V$, producer surplus $v_{L}$, and type-H consumers surplus $\mu_{0}\Delta v$:
$V+v_{L}+\mu_{0}\Delta v.$
As $\pi\left(l|L\right)=\pi\left(l|H\right)=1$ is optimal when $\mu_{0}<v_{L}/v_{H}$ from Proposition 2, it follows
Corollary 2.
Suppose $\mu_{0}<v_{L}/v_{H}$. The no-information data trade is always optimal.
@@ 4.2 Intermediary Optimality for $\mu_{0}<v_{L}/v_{H}$
The optimal data trade can disclose no information about consumers. Since the product market has included the type-L consumers when $\mu_{0}<\frac{v_{L}}{v_{H}}$, there is no space for the producer to extract type-L consumer surplus from the extra information of data trade. Hence, there is no price discrimination in the product market.
4.3 Optimal Data Trade
To be clear on the tradeoff of data trade, we summarize the results on data trade in Table 1. Several implications follow the optimal data trade results.
Table 1: Optimal Data Trade
Prior Mismatch Data Trade
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[0,C\right]$
$\delta^{2}<\Delta v$ $\pi\left(l|L\right)=\frac{\delta^{2}}{\Delta v},\pi\left(l|H\right)\in\left[0,C\frac{\delta^{2}}{\Delta v}\right]$
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[0,1\right]$
$\delta^{2}<\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[1-\frac{\delta^{2}}{\Delta v},1\right]$
@@ 4.3 Optimal Data Trade
First, the intermediary will not sell full information in the optimal mechanism when $\delta^{2}<\Delta v$222222 The result is mainly from the cross-market effect of data, different from the ratchet effect. To see this, take the case of $\mu_{0}>\frac{v_{L}}{v_{H}}$ as an example. In the optimal mechanism, the service profit is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and the data trade fee is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. In contrast, if the data trade is full information, the service profit decreases to $V+\mu_{0}\left(\delta^{2}-\Delta v\right)$, while the data trade fee increases to $\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. The loss in service profit of the amount $\mu_{0}\pi\left(l|H\right)\Delta v-\mu_{0}\left(\delta^{2}-\Delta v\right)$ is larger than the amount of increment in the data profit $\mu_{0}\pi\left(l|H\right)\Delta v$ when $\delta^{2}<\Delta v$. To preserve the service profit, the intermediary would prefer to sell only partial or no information. [^23] The total payoff from the full information $V+\mu_{0}\left(\delta^{2}-\Delta v\right)+\left(1-\mu_{0}\right)v_{L}$ is smaller than the one from the optimal data trade $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta}$ when $\delta^{2}<\Delta v$. Instead, when $\delta^{2}\geq\Delta v$ selling full information would be optimal for the intermediary. The following result directly follows.
Corollary 3.
When $\delta^{2}<\Delta v$, the full information data trade cannot be optimal.
Second, the privacy protection laws should be careful about the suit levels of protection depending on the market environment. The privacy protection laws usually work as the intermediary’s commitment on how much data to trade. The highest privacy protection is sometimes optimal (e.g. small $\mu_{0}$), while trading full data is sometimes optimal (e.g., large $\delta^{2}$).
4.4 Welfare and Privacy Protection
The baseline framework with a holistic viewpoint provides an opportunity to evaluate the effects of the laws of privacy protection comprehensively: Taking into account both data collecting and selling. The booming digital economy featuring data trade and transfer brings wide concerns about privacy protections, which attracts attention from both the public and regulatory authorities. Several laws have been established across different countries. Despite the importance of privacy protections, a comprehensive understanding of the potential effects of regulations on the data market is still missing.
@@ 4.4 Welfare and Privacy Protection
When the laws ban the data trade to pursue the highest level of consumer privacy protection, the service market, and product market work separately because they are not linked by the intermediary’s data trade. The data shared with the intermediary will no longer be used in another market against the consumers’ interests. In this environment, the type-L consumers’ total surplus is zero. The type-H consumers get some surplus when the product price is $v_{L}$, which corresponds to the case of $\mu_{0}<\frac{v_{L}}{v_{H}}$. In the product market, the producer extracts all the surplus under the prior $\mu_{0}$. The producer’s surplus is $R_{0}=\max\left\{v_{L},v_{H}\mu_{0}\right\}$.
Proposition 3.
@@
Suppose there exists no data trade.
@@
    In the product market, the producer sets price $p_{0}$. The type-L consumers surplus is 00. The type-H consumers surplus is $\Delta v$ when $\mu_{0}<\frac{v_{L}}{v_{H}}$ and is 00 when $\mu_{0}>\frac{v_{L}}{v_{H}}$. The producer gets $\max\left\{v_{L},\mu_{0}v_{H}\right\}$.
Without data trade, social welfare is $V+\mu_{0}v_{H}$ when $\mu_{0}>\frac{v_{L}}{v_{H}}$, and $V+v_{L}+\mu_{0}\Delta v$ when $\mu_{0}<\frac{v_{L}}{v_{H}}$.
To investigate the effects of banning data trade, we compare the surplus distribution and welfare with and without the data market.
@@
With Data Market Without Data Market Prior Mismatch Intermediary Producer Consumer $H$ Consumer $L$ Intermediary Producer Consumer $H$ Consumer $L$ $\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}$ $\mu_{0}v_{H}$ 0 0 $V$ $\mu_{0}v_{H}$ 0 0 $\delta^{2}<\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}\Delta v$ $\mu_{0}v_{H}$ 0 0 $V$ $\mu_{0}v_{H}$ 0 0 $\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V$ $v_{L}$ $\Delta v$ 00 $V$ $v_{L}$ $\Delta v$ 00 $\delta^{2}<\Delta v$ $V$ $v_{L}$ $\Delta v$ 00 $V$ $v_{L}$ $\Delta v$ 00
Table 3: Surplus Change Led by Data Trade Ban
Prior Mismatch Intermediary Producer Consumer $H$ Consumer $L$
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$downarrow -- -- --
$\delta^{2}<\Delta v$downarrow -- -- --
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$- -- -- --
$\delta^{2}<\Delta v$- -- -- --
Surplus.
In Table 2, we summarize the consumers, the producer, and the intermediary surplus with and without data trade ban. The data trade ban does not affect the producer surplus and consumers surplus. The producer surplus is always the producer’s revenue in the product market without buying data, $R_{0}$. This is because the intermediary has the full bargaining power over the data trade surplus and the producer does not get any share of the data trade surplus even when there is a data market. The type-L consumers surplus is always zero: The intermediary extracts their surplus in the service market and they don’t have any surplus in the product market. The type-H consumers surplus is $\Delta v$ in the product market.
Banning data trade affects the intermediary only when $\mu_{0}>\frac{v_{L}}{v_{H}}$. The intermediary can only extract the service surplus $V$ without the data market. However, the intermediary gets better off by serving an extra proportion of type-L consumers due to the positive market inclusion effect of data trade. With data trade, the producer is able to price discriminate and extract more consumers surplus by $\left(1-\mu_{0}\right)v_{L}\min\left\{\Delta v,1\right\}$. Changes in surplus are summarized in Table 3.
Social Welfare.

### v3

@@ 4.1 Special Case of Binary Service
When the horizontal differentiation level is below the vertical one, $\delta^{2}<\Delta v$, the mismatch cost in service is low as a penalty for the type-H consumers’ misreport. The mismatch itself cannot provide sufficient incentive for the type-H consumers to reveal their information. Instead, the intermediary would like to lower $\pi\left(l|L\right)$ to incentivize truth-telling to collect information. In this case, some type-L consumers are excluded from the product market.
4.2 Intermediary Optimal Mechanisms
We return to the complete characterization of optimal data collection and data trade under the service space $\mathcal{X}=\mathbb{R}$. The analysis shows that the fundamental tradeoff of data trade works under the general service, while admitting flexible service distortions and their effects on consumers’ incentives of sharing data.
When $\mu_{0}<v_{L}/v_{H}$, even allowing general service provisions, there is no distortion of service in the intermediary optimal mechanism. The results on service fee, data trade, and data fee are the same as those in 1. We state them as follows.
Proposition 2 (Optimal design—small $\mu_{0}$).
Suppose $\mu_{0}<v_{L}/v_{H}$. In the optimal mechanism, there is no distortion in service provisions $x_{L}=L$, $x_{H}=H$, the service fees are $f_{L}=V$, $f_{H}=V+[\pi(l|H)-1]\Delta v$, and the intermediary’s data trade is
$\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\Big[\max\Big\{0,1-\frac{\delta^{2}}{\Delta v}\Big\},1\Big],$
and the data trade fee is $T=\mu_{0}[1-\pi\left(l|H\right)]\Delta v$. In this case, social welfare is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}.$
Now we consider the case of $\mu_{0}>v_{L}/v_{H}$. Although there is no distortion in service provisions (and the service fees, data trade are as in part 1 of 1) for $\delta^{2}\geq\Delta v$, the service provided to the type-L is downward distorted for $\delta^{2}<\Delta v$. We summarize the results for $\mu_{0}>v_{L}/v_{H}$ in the following proposition.
Proposition 3 (Optimal design—large $\mu_{0}$).
Suppose $\mu_{0}>v_{L}/v_{H}$, in the optimal mechanisms:
  1. 1.
    For $\frac{\delta^{2}}{\Delta v}\geq 1$, there are no distortion in service provisions $x_{L}=L$, $x_{H}=H$, the service fees are $f_{L}=V$, $f_{H}=V+\pi(l|H)\Delta v$, and the intermediary’s data trade is
    and the data trade fee is $T=-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}$, and the intermediary’s total revenue is $V+(1-\mu_{0})v_{L}$. In this case, social welfare is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}$.
  2. 2.
    For $\delta^{2}<\Delta v$, there is a downward distortion in service provided to type-$L$ consumers, but there is no service distortion for type-$H$ consumers $x_{H}=H$. The following holds:
    1. (a)
      $\pi(l|L)=1,\text{ }\pi(l|H)\in[0,C]$
      the data trade fee is $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}$ and the intermediary’s total revenue is $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}$. In this case, social welfare is $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$.
    2. (b)
      If $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{L}+v_{H}}$, then the service for type-$L$ is $x_{L}=L-\frac{\delta}{\Delta v}v_{L}$, the service fees are $f_{L}=V-(\frac{\delta}{\Delta v}v_{L})^{2}$, $f_{H}=V+\pi(l|H)\Delta v$; and the data trade is
      $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v},\text{ }\pi(l|H)\in\left[0,C\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}\right]$
      the data trade fee is $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}\frac{\delta^{2}}{(\Delta v)^{2}}(v_{L}+v_{H})$, and the intermediary’s total revenue is $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}$. In this case, social welfare is $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\mu_{0}v_{H}$.
When $\mu_{0}>v_{L}/v_{H}$ and the horizontal differentiation level is high relative to the vertical one (i.e., $\delta^{2}\geq\Delta v$), there is no distortion of service provisions to consumers. This is because the mismatch cost in service is so high that consumers are willing to report the private information without any further distortion in service level. The intermediary sets the highest $\pi\left(l|L\right)=1$ and does not need to distort any service level to maintain the type-H’s incentive.
However, when $\mu_{0}>v_{L}/v_{H}$ and the horizontal differentiation level is below the vertical one, $\delta^{2}<\Delta v$, the intermediary needs to downward distort the type-L’s service. In this case, the penalty of the wrong service cannot provide enough incentive. The intermediary will manipulate the type-H consumers’ incentive via the other two instruments: the probability of surplus in the product and service provision. Although distorting the service will reduce the surplus extracted from the type-Ls in the service market, the intermediary can extract more surplus from the type-H consumers in the product market. When the benefit of more surplus from type-H consumers outweighs the cost of reduced surplus from type-L consumers, the intermediary would prefer to distort the service to the type-L consumers. This happens when the population of type-H consumers is large enough $\mu_{0}>v_{L}/v_{H}$.
Table 1: Main Results of Optimal Mechanism
Prior Condition Service Provision Data Trade $\mu_{0}>\frac{v_{L}}{v_{H}}$ $\frac{\delta^{2}}{\Delta v}\geq 1$ no distortion $\pi(l|L)=1,\pi(l|H)\in[0,C]$ $\frac{\Delta v}{v_{H}+v_{L}}\leq\frac{\delta^{2}}{\Delta v}<1$ downward distortion at bottom $\pi(l|L)=1,\pi(l|H)\in[0,C]$ $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{H}+v_{L}}$ downward distortion at bottom $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v},\pi(l|H)\in[0,C\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}]$ $\mu_{0}<\frac{v_{L}}{v_{H}}$ $\frac{\delta^{2}}{\Delta v}\geq 1$ no distortion $\pi(l|L)=1,\pi(l|H)\in[0,1]$ $\frac{\delta^{2}}{\Delta v}<1$ no distortion $\pi(l|L)=1,\pi(l|H)\in[1-\frac{\delta^{2}}{\Delta v},1]$
Following 2 and 3 (summarized in Table 1) directly, we can discuss some special data trade policies, such as full privacy and full disclosure.
Corollary 1 (Full privacy).
If $\mu_{0}>v_{L}/v_{H}$, the null information data trade cannot be optimal. If $\mu_{0}<v_{L}/v_{H}$, the null information data trade is always optimal.
@@ 4.2 Intermediary Optimal Mechanisms
It says that full privacy cannot be optimal for $\mu_{0}>v_{L}/v_{H}$. [^21] Calzolari and Pavan (2006b) offer conditions for full privacy to be optimal. Their condition about the same sign of single-crossing conditions is violated in our setup. To see it by contradiction. Suppose in the optimal mechanism, $\pi\left(l|H\right)=\pi\left(l|L\right)$. Then the obedience constraint (16) requires $(1-\frac{v_{L}}{v_{H}})\mu_{0}\leq\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)$ which contradicts with $\mu_{0}>\frac{v_{L}}{v_{H}}$. Intuitively, when $\mu_{0}>\frac{v_{L}}{v_{H}}$ any extra information will make it easier for the producer to price discriminate the consumers. This is because in the case of $\mu_{0}>\frac{v_{L}}{v_{H}}$ the market inclusion effect is positive: without any data trade all the type-L consumers are excluded in the product market.
@@
But the optimal data trade can feature full privacy for $\mu_{0}<v_{L}/v_{H}$. As the product market has included the type-L consumers when $\mu_{0}<v_{L}/v_{H}$, there is no space for the producer to extract type-L consumer surplus from the extra information of data trade.
Corollary 2 (Full disclosure).
For $\mu_{0}>v_{L}/v_{H}$, selling full information is optimal to the intermediary if and only if $\frac{\delta^{2}}{\Delta v}\geq\frac{\Delta v}{v_{L}+v_{H}}$. For $\mu_{0}<v_{L}/v_{H}$, selling full information is optimal to the intermediary if and only if $\frac{\delta^{2}}{\Delta v}\geq 1$.
@@
For instance, in the case of $\mu_{0}>v_{L}/v_{H}$, the intermediary will not sell full information in the optimal mechanism when $\delta^{2}<(\Delta v)^{2}/(v_{L}+v_{H})$. To see this by contradiction, in the optimal mechanism, the service profit is $V+\mu_{0}\pi(l|H)\Delta v-(1-\mu_{0})(\frac{\delta}{\Delta v}v_{L})^{2}$ and the data profit is $\mu_{0}\pi\left(l|H\right)\Delta v+(1-\mu_{0})v_{L}\frac{\delta^{2}}{\Delta v}\frac{v_{L}+v_{H}}{\Delta v}$. Suppose the data policy sells full information, the service profit decreases to $V-(1-\mu_{0})(\frac{\Delta v-\delta^{2}}{2\delta})^{2}$, while the data profit increases to $(1-\mu_{0})v_{L}$. However, the amount of loss in service profit is larger than that of increment in the data profit when $\delta^{2}<(\Delta v)^{2}/(v_{L}+v_{H})$. [^22] It is because given $\delta^{2}<(\Delta v)^{2}/(v_{L}+v_{H})$, $V+\mu_{0}\pi(l|H)\Delta v-(1-\mu_{0})(\frac{\delta}{\Delta v}v_{L})^{2}-\left(V+(1-\mu_{0})(\frac{\Delta v-\delta^{2}}{2\delta})^{2}\right)$ $\hskip 72.26999pt>(1-\mu_{0})v_{L}-\left(-\mu_{0}\pi\left(l|H\right)\Delta v+(1-\mu_{0})v_{L}\frac{\delta^{2}}{\Delta v}\frac{v_{L}+v_{H}}{\Delta v}\right).$ To preserve the service profit, the intermediary would prefer to sell partial information. Instead, when $\delta^{2}\geq(\Delta v)^{2}/(v_{L}+v_{H})$, selling full information can be optimal for the intermediary, even considering the cross-market effect of data trade.
Another interesting observation about the service fees by contrasting the results in 2 with 3 is that the high type consumers end up paying more service fee in the data-sourcing market for a larger $\mu_{0}$; but the low type consumers pay less (equal) service fee if $\delta^{2}<\Delta v$ ($\delta^{2}\geq\Delta v$ ) for a larger $\mu_{0}$.
4.3 Regulation of Data Sharing
The above framework with a holistic viewpoint, considering data sourcing, selling, and using, provides an opportunity to investigate the data market regulations in an integrated way. The regulatory policies have been established across different countries to address concerns about consumer privacy, led by data transfers across markets. However, a comprehensive understanding of the potential effects of regulations on the data market is still missing. The subsection studies the privacy regulation (or the regulation of data sharing). [^23] Section 5 discusses the intermediary’s partial bargaining power and its regulation.
Efficiency.
In the socially optimal allocation, each consumer is provided a correct service and offered a product, and hence the socially optimal welfare is $V+\mu_{0}v_{H}+(1-\mu_{0})v_{L}$. Following the propositions 2 and 3 (summarized in Table 1), it is trivial that the intermediary optimal allocation is efficient for $\mu_{0}<v_{L}/v_{H}$. However, for $\mu_{0}>v_{L}/v_{H}$, the optimal mechanism of the intermediary is efficient if and only if $\delta^{2}\geq\Delta v$.
Lemma 3 (Efficient mechanism).
Suppose $\mu_{0}>v_{L}/v_{H}$. The intermediary’s optimal mechanisms are efficient if and only if $\delta^{2}\geq\Delta v$.
When $\delta^{2}\geq\Delta v$, there is no distortion of service provisions, and the optimal data trade which fully discloses information about consumers’ willingness to pay induces perfect price discrimination in the product market. Hence, the intermediary optimal mechanism is efficient under certain market differentiation conditions. The efficiency result implies that the data market regulations that intervene in the amount of data may lead to inefficiency.
Surplus and Welfare.
@@ 4.3 Regulation of Data Sharing
To investigate the effects of banning data trade, we compare the surpluses and welfare with and without the data market. When the laws ban the data trade to pursue the highest level of consumer privacy protection, the service market and product market work separately because they are not linked by the intermediary’s data trade. The data shared with the intermediary will no longer be used in another market against the consumers’ interests. In this environment, the type-L consumers’ total surplus is zero. The type-H consumers get some surplus when the product price is $v_{L}$, which corresponds to the case of $\mu_{0}<\frac{v_{L}}{v_{H}}$. In the product market, the producer extracts all the surplus under the prior $\mu_{0}$. The producer’s surplus is $R_{0}=\max\left\{v_{L},v_{H}\mu_{0}\right\}$.
Lemma 4 (Results without data market).
@@
Suppose there exists no data trade. Social welfare is $V+\mu_{0}v_{H}$ for $\mu_{0}>v_{L}/v_{H}$, and is $V+v_{L}+\mu_{0}\Delta v$ for $\mu_{0}<v_{L}/v_{H}$.
@@
    In the product market, the producer sets price $p_{0}$. The type-L consumers surplus is 0. The type-H consumers surplus is $\Delta v$ when $\mu_{0}<v_{L}/v_{H}$ and is 0 when $\mu_{0}>v_{L}/v_{H}$. The producer gets $\max\left\{v_{L},\mu_{0}v_{H}\right\}$.
Table 2 summarizes the consumers, the producer, and the intermediary surplus with and without data trade ban. The data trade ban does not affect the producer surplus and consumers surplus. The producer surplus is always the producer’s revenue in the product market without buying data, $R_{0}$. This is because the intermediary has the full bargaining power over the data trade surplus and the producer does not get any share of the data trade surplus even when there is a data market. The type-L consumers surplus is always zero: The intermediary extracts their surplus in the service market and they don’t have any surplus in the product market. The type-H consumers surplus is $\mu_{0}\Delta v$ in the product market.
Banning data trade affects the intermediary’s payoff only when $\mu_{0}>v_{L}/v_{H}$. Note that selling no information is optimal to the intermediary for $\mu_{0}<v_{L}/v_{H}$ by 2, hence, banning data trade won’t hurt the intermediary for $\mu_{0}<v_{L}/v_{H}$. However, it will hurt the intermediary for $\mu_{0}>v_{L}/v_{H}$, because selling no information is not optimal from 3. The intermediary can only extract the service surplus $V$ without the data market, but gets better off by serving an extra proportion of type-L consumers with the data market.
@@
With Data Market Without Data Market Primitive Condition Intermediary Producer Consumers Intermediary Producer Consumers $\mu_{0}>\frac{v_{L}}{v_{H}}$ $\frac{\delta^{2}}{\Delta v}\geq 1$ $V+\left(1-\mu_{0}\right)v_{L}$ $\mu_{0}v_{H}$ 0 $V$ $\mu_{0}v_{H}$ 0 $\frac{\Delta v}{v_{H}+v_{L}}\leq\frac{\delta^{2}}{\Delta v}<1$ $V+(1-\mu_{0})v_{L}-(1-\mu_{0})(\frac{\Delta v-\delta^{2}}{2\delta})^{2}$ $\mu_{0}v_{H}$ 0 $V$ $\mu_{0}v_{H}$ 0 $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{H}+v_{L}}$ $V+(1-\mu_{0})v_{L}v_{H}(\frac{\delta}{\Delta v})^{2}$ $\mu_{0}v_{H}$ 0 $V$ $\mu_{0}v_{H}$ 0 $\mu_{0}<\frac{v_{L}}{v_{H}}$ $\frac{\delta^{2}}{\Delta v}\geq 1$ $V$ $v_{L}$ $\mu_{0}\Delta v$ $V$ $v_{L}$ $\mu_{0}\Delta v$ $\frac{\delta^{2}}{\Delta v}<1$ $V$ $v_{L}$ $\mu_{0}\Delta v$ $V$ $v_{L}$ $\mu_{0}\Delta v$